OpenStudy (anonymous):

I need to write the radical expression in exponential form. 8 divided by 7th. root of x to the 15th. power. I think it is8x^7/15. Is that correct?

OpenStudy (anonymous):

Almost.

OpenStudy (anonymous):

\[\large \frac{8}{\sqrt[7]{x^{15}}} \rightarrow 8 \cdot x^{-15/7}\]

OpenStudy (anonymous):

\[\large 8 \cdot x^{7/15} = 8 \cdot \sqrt[15]{x^7}\]

OpenStudy (anonymous):

I thought the denominator would be multiplied by x ^7/15 and also the numerator by the same

OpenStudy (anonymous):

Hmm, like a rationalizing the denominator trick?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You need to multiply by something else. x^(15/7) × x^(7/15) = x^(274/105)

OpenStudy (anonymous):

what happened to the denominaif it wasn't rationalized

OpenStudy (anonymous):

Do you want your answer with no negative exponents? And/or no radicals in the denominator?

OpenStudy (anonymous):

not sure. My choices for answers doesn't include any denominators. your answer is one of my choices but what happened to the denominator?

OpenStudy (anonymous):

I expressed it using a negative exponent.

OpenStudy (anonymous):

got it ok thanks

OpenStudy (anonymous):

1/x = x^-1, etc.

OpenStudy (anonymous):

yep